OpenStudy (anonymous):

How do I work out the equation, -5(x – 4) + 2 = -13 ? Help!

mathslover (mathslover):

ok : \(\huge{a(b+c)=ab+ac}\)

OpenStudy (anonymous):

Use the distributive property

mathslover (mathslover):

similarly \(\Huge{-5(x-4)=-5(x)-5(-4)}\)

mathslover (mathslover):

\[\Huge{-5x+20+2=-13}\] \[\Huge{-5x+22=-13}\] \[\Huge{-5x=-35}\] \[\Huge{x=\frac{-35}{-5}}\] \[\Huge{x=7}\]

OpenStudy (anonymous):

how do you get the numbers so big?

OpenStudy (anonymous):

ohh, thank you !